# A bound-preserving Runge–Kutta discontinuous Galerkin method with compact stencils for hyperbolic conservation laws

### v1

@@ 2.2.1 Quadrature and convex decomposition of cell averages
Here, $\omega^{\mathrm{B}}_{\beta,\ell,K}$ and $x_{\beta,\ell,K}$ ($\beta=1,\cdots,{s}_{\ell}$) are the associated quadrature weights and abscissae. This quadrature rule should be exact for the integration of constants, i.e., satisfies the Proposition 2.1.
@@
An illustration of the quadratures that enjoy Proposition 2.1 and Proposition 2.2 and are used in third-order and fourth-order schemes on 2D structured meshes is given in Figure 1. Note that in particular, when taking $v(x)\equiv 1$, (2.14) in Proposition 2.2 implies
@@ 2.2.3 Hyperbolic systems: compressible Euler equations
$m=\rho w,\quad E=\frac{|m|^{2}}{2\rho}+pe,\quad\text{and}\quad p=(\gamma-1)\rho e.$ (2.23)
@@
In practice, $\epsilon$ can be set to $10^{-8}$ or $10^{-13}$ in the computation [38, Section 4.2]. Define the numerical admissible set for the compressible Euler equations:
@@ 3.1.2 Hyperbolic systems: compressible Euler equations
For Euler equations in (2.22) with the invariant region $B$ defined in (2.24), if $\lambda a_{0}\leq c_{\mathcal{G}}$, then for any unit vector $\nu$, we have
@@
Recall that $p=(\gamma-1)(E-\frac{1}{2}\frac{|m|^{2}}{\rho})$. Let the bracket []delimited-[][$\cdot$][ ⋅ ] below denote the dependence. By performing a direct calculation similar to the one in [40, Remark 2.4], we have
@@ 3.2 Bound-preserving framework for cRKDG schemes
Consider the $i$th stage of a cRKDG method $u_{h}^{(i)}$. Suppose $u_{h}^{(i)}$ can be written as a convex combination of forward-Euler steps with $\mathcal{F}$ and ±$\pm\mathcal{G}$ for previous stages. In other words, it can be written as a convex combintation of
$u_{h}^{(j)}+\Delta t\mu_{ij}^{\mathcal{F}}\mathcal{F}\left(u_{h}^{(j)}\right)\quad\text{and}\quad u_{h}^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}\left(u_{h}^{(j)}\right)\quad\text{with}\quad j<i,\quad\mu_{ij}^{\mathcal{F}},\mu_{ij}^{\pm\mathcal{G}}>0.$ (3.21)
@@
Note that under the given time step constraint (3.22), from Propositions 2.3, 3.1, 2.5 and 3.3, we have
@@
Consider a cRKDG method. Suppose each of its stage can be written as a convex of forward-Euler steps with $\mathcal{F}$ and ±$\pm\mathcal{G}$ for previous stages. Under the CFL condition
@@
For relativistic hydrodynamics, the weak bound-preserving property for $\mathcal{F}$ has been proved in [24] for Cartesian meshes; the proof of the Lax–Friedrichs splitting property can be found at [32, Lemma 2.3] and also in [24].
@@ 3.3.1 The second-order case
Then, we minimize the coefficients $\mu_{ij}^{\mathcal{F}}$ and $\mu_{ij}^{\pm\mathcal{G}}$. With the order conditions and the assumption that all parameters are nonnegative, this leads to the following optimization problem:
@@
We use the $\mathtt{fminimax}$ function in MATLAB to solve the minimax constraint problem. The optimal solution is non-unique. A simple one we found is
@@ 4.2 Convergence study
For the tested cRKDG algorithms, we combine $\mathds{P}^{k}$- or $\mathds{Q}^{k}$-DG schemes for $k=1,2,3$ with $(k+1)$th-order RK time discretization, as specified in (3.36), (3.45), and (3.52), respectively. Let $\mathtt{err}_{\Delta x}$ denote the error on a mesh with resolution $\Delta x$. The convergence rate is computed by $\ln(\mathtt{err}_{\Delta x}/\mathtt{err}_{\Delta x/2})/\ln(2)$.

### v2

@@ 2.2.1 Quadrature and convex decomposition of cell averages
Here, $\omega^{\mathrm{B}}_{\beta,\ell,K}$ and $x_{\beta,\ell,K}$ ($\beta=1,\cdots,{s}_{\ell}$) denote the associated quadrature weights and abscissae. This quadrature rule should be exact for the integration of constants, i.e., it satisfies the Proposition 2.1.
@@
An illustration of the quadratures that enjoy Propositions 2.1 and 2.2 and are used in third-order and fourth-order schemes on 2D structured meshes is given in Figure 1. Note that in particular, when taking $v(x)\equiv 1$, (2.14) in Proposition 2.2 implies
@@ 2.2.2 Scalar hyperbolic conservation laws
Here $c_{\mathrm{\mathcal{F}}}$ is a positive constant. In the case that the same quadrature rule is used along all edges, $c_{\mathrm{\mathcal{F}}}=\min_{\beta,\ell,K}\frac{\omega_{\beta,\ell,K}^{\mathrm{I}}}{\omega_{\beta,\ell,K}^{\mathrm{B}}}$ for Cartesian meshes and $c_{\mathrm{\mathcal{F}}}=\min_{\beta,\ell,K}\frac{|e_{\ell,K}|}{\sum_{l=1}^{3}|e_{l,K}|}\frac{\omega_{\beta,\ell,K}^{\mathrm{I}}}{\omega_{\beta,\ell,K}^{\mathrm{B}}}$ for triangular meshes.
@@ 2.2.3 Hyperbolic systems: compressible Euler equations
$m=\rho w,\quad E=\frac{|m|^{2}}{2\rho}+\rho e,\quad\text{and}\quad p=(\gamma-1)\rho e.$ (2.23)
@@
Here $c_{\mathrm{\mathcal{F}}}$ is a positive constant. In the case that the same quadrature rule is used along all edges, $c_{\mathrm{\mathcal{F}}}=\min_{\beta,\ell,K}\frac{\omega_{\beta,\ell,K}^{\mathrm{I}}}{\omega_{\beta,\ell,K}^{\mathrm{B}}}$ for Cartesian meshes and $c_{\mathrm{\mathcal{F}}}=\min_{\beta,\ell,K}\frac{|e_{\ell,K}|}{\sum_{l=1}^{3}|e_{l,K}|}\frac{\omega_{\beta,\ell,K}^{\mathrm{I}}}{\omega_{\beta,\ell,K}^{\mathrm{B}}}$ for triangular meshes.
@@
In practice, $\epsilon$ can be set to $10^{-8}$ or $10^{-13}$ in the computation [39, Section 4.2]. Define the numerical admissible set for the compressible Euler equations:
@@ 3.1.2 Hyperbolic systems: compressible Euler equations
For Euler equations in (2.22), with the invariant region $B$ defined in (2.24), and for any unit vector $\nu$, we have
@@
Recall that $p=(\gamma-1)(E-\frac{1}{2}\frac{|m|^{2}}{\rho})$. Let the bracket []delimited-[][$\cdot$][ ⋅ ] below denote the dependence. By performing a direct calculation similar to the one in [41, Remark 2.4], we have
@@
Remark 2.
A simple alternative proof of Proposition 3.2 can be obtained using the geometric quasilinearization. We refer to Appendix A for details.
@@ 3.2 Bound-preserving framework for cRKDG schemes
Consider the $i$th stage of a cRKDG method $u_{h}^{(i)}$. Suppose that the cell average of $u_{h}^{(i)}$ can be written as a convex combination of those of forward-Euler steps with $\mathcal{F}$ and ±$\pm\mathcal{G}$ for previous stages. In other words, it can be written as a convex combintation of
$\overline{u_{h}^{(j)}+\Delta t\mu_{ij}^{\mathcal{F}}\mathcal{F}\left(u_{h}^{(j)}\right)}\quad\text{and}\quad\overline{u_{h}^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}\left(u_{h}^{(j)}\right)}\quad\text{with}\quad j<i,\quad\mu_{ij}^{\mathcal{F}},\mu_{ij}^{\pm\mathcal{G}}>0.$ (3.21)
@@
Note that under the given time step constraint (3.22), from Propositions 2.3, 3.1, 2.5, and 3.3, we have
@@
With Lemma 3.4, we can prove the main result of the paper, Theorem 3.5, whose detailed proof is provided in Appendix B.
@@
Consider a cRKDG method. Suppose that, without using limiters, each of its stages can be written as a convex combination of forward-Euler steps involving $\mathcal{F}$ and ±$\pm\mathcal{G}$ for previous stages. Under the CFL condition
@@
Remark 3 (CFL condition for stability).
In the analysis of this paper, the time step constraint is imposed solely to ensure the weak bound-preserving property of the cRKDG method. In practice, however, the time step must also satisfy the stability requirement. The CFL numbers for the $(k+1)$th-order cRKDG methods with $\mathds{P}^{k}$ elements, where $k=1,2,3$, are 0.333, 0.178, and 0.103, respectively [1]. These values are obtained via a Fourier-type analysis. For comparison, the corresponding CFL numbers ensuring stability for the RKDG methods are 0.333, 0.209, and 0.145, respectively [9].
Remark 4 (Applying limiter after each stage).
We clarify that, the limiter is applied after each cRKDG stage, not after each forward-Euler step. The decomposition of an RK stage into a convex combination of forward-Euler steps is used solely to prove the weak bound-preserving property of each stage and is not required in the implementation of the solver.
@@
For relativistic hydrodynamics, the weak bound-preserving property for $\mathcal{F}$ has been proved in [25] for Cartesian meshes; the proof of the Lax–Friedrichs splitting property can be found at [33, Lemma 2.3] and also in [25].
@@ 3.3.1 The second-order case
Then, we minimize the coefficients $\mu_{ij}^{\mathcal{F}}$ and $\mu_{ij}^{\pm\mathcal{G}}$. There are three types of constraints that must be considered. First, (3.31) must be a convex combination of forward-Euler steps, which requires $0\leq\alpha,\beta\leq 1$. Second, the forward-Euler step with the DG operator $\mathcal{F}$ must advance the solution forward in time; in addition to the previous condition, this requires $0\leq b_{1},b_{2}$. Finally, to achieve second-order temporal accuracy, the RK method must satisfy the order conditions $b_{1}+b_{2}=1$ and $b_{2}a_{21}=\frac{1}{2}$, which can be derived by matching the coefficients in the Taylor expansion of the discrete solution with those of the continuous solution. For further details, we refer to [20, Section II].
With the aforementioned constraints, we have the following optimization problem:
@@
We use the fminimax function in MATLAB to solve the minimax constraint problem. The optimal solution is non-unique. A simple one we found is
@@
Remark 10.
The key to the parameter search of the cRKDG method lies in expressing it in a generalized Shu–Osher form, as in (3.29a) and (3.31). We consider it “generalized” because we distinguish the forward-Euler steps that share the same stage $u^{(i)}$ but involve different spatial operators. Compared with the standard SSP-RKDG method, the parameter search of the cRKDG method has the following constraint and flexibility.
  1. 1.
    (Constraint) When written in Butcher form, the final stage of the cRKDG method must not contain the operator $\mathcal{G}$. This restriction is equivalent to imposing additional constraints on the generalized Shu–Osher coefficients.
  2. 2.
    (Flexibility) When expressed in the generalized Shu–Osher form, the cRKDG method allows the coefficients in front of $\mathcal{G}$ to be negative.
The constraint tends to reduce the SSP coefficients of the cRKDG method relative to those of the standard SSP-RKDG method, resulting in a more restrictive time-step condition. On the other hand, the flexibility enables the method to overcome certain order barriers, as will be demonstrated in the fourth-order case. This flexibility is akin to allowing downwind stages in the standard SSP-RKDG scheme, which can also overcome the order barrier but may introduce complications when imposing downwind flux at the outflow boundary. However, thanks to the compactness of $\mathcal{G}$, the cRKDG method does not face this issue.
@@ 4.2 Convergence study
For the tested cRKDG algorithms, we combine $\mathds{P}^{k}$- or $\mathds{Q}^{k}$-DG schemes for $k=1,2,3$ with $(k+1)$th-order RK time discretization, as specified in (3.36), (3.45), and (3.52), respectively. Let $\texttt{err}_{\Delta x}$ denote the error on a mesh with resolution $\Delta x$. The convergence rate is computed by $\ln(\texttt{err}_{\Delta x}/\texttt{err}_{\Delta x/2})/\ln(2)$.
@@ 5 Conclusions
Acknowledgment
We would like to thank the two reviewers for their valuable comments, which have helped improve the paper.
Appendix A An alernative proof of Propsoition 3.2
In [32, Theorem 5.1], Wu and Shu proved that the invariant region for the Euler equations (2.22) and (2.23) can be equivalently formulated as
$B=\{u=(\rho,m^{\top},E)^{\top}:u\cdot n_{1}>0,u\cdot n(w_{*})>0\ \forall w_{*}\in\mathbb{R}^{d}\}.$ (A.1)
Here $n_{1}=(1,0^{\top},0)^{\top}$ and $n(w_{*})=(\frac{w_{*}^{2}}{2},-w_{*},1)^{\top}$. Note for any $u\in B$, we have
(±$(\pm a_{0}^{-1}f(u)\cdot v)\cdot n_{1}=\pm\frac{w\cdot v}{a_{0}}u\cdot n_{1}<u\cdot n_{1}.$ (A.2)
Moreover, with the arithmetic-geometric inequality and the fact $|v|=1$, we have
(±$(\pm a_{0}^{-1}f(u)\cdot v)\cdot n(w_{*})=$$\pm\frac{w\cdot v}{a_{0}}u\cdot n(w_{*})\pm\frac{pv\cdot(w-w_{*})}{a_{0}}$ (A.3)
$\leq$ $\frac{|w\cdot v|}{a_{0}}u\cdot n(w_{*})+\frac{\sqrt{\frac{\gamma-1}{2}\frac{p}{\rho}}}{a_{0}}\left(\frac{p}{\gamma-1}+\frac{1}{2}\rho|w-w_{*}|^{2}\right)$
$=$ $\frac{|w\cdot v|+\sqrt{\frac{\gamma-1}{2}\frac{p}{\rho}}}{a_{0}}u\cdot n(w_{*})<u\cdot n(w_{*}).$
As a result, (A.2) implies that $(u\pm a_{0}^{-1}f(u)\cdot v)\cdot n_{1}>0$, and (A.3) implies that $(u\pm a_{0}^{-1}f(u)\cdot v)\cdot n(w_{*})>0$. Therefore, the two conditions in (A.1) are satisfied, which gives $u\pm a_{0}^{-1}f(u)\cdot v\in B$.
Appendix B Proof of Theorem 3.5
In this section, we omit the subscripts $h$ and $K$ for simplicity.
B.1 Main technicality: the second-order case
The main technical challenge in the proof arises from applying the limiter after each RK stage. Before presenting the rigorous proof, we provide a brief explanation using the second-order bound-preserving cRKDG method as an example.
If no limiter is applied, Scheme 1 in (B.1) and Scheme 2 in (B.2) are equivalent.
Scheme 1:
$u^{(2)}$ $=u^{n}+\frac{\Delta t}{2}\mathcal{G}(u^{n}),$ (B.1a)
$u^{n+1}$ $=u^{n}+\Delta t\mathcal{F}(u^{(2)}).$ (B.1b)
Scheme 2:
$u^{(2)}$ $=u^{n}+\frac{\Delta t}{2}\mathcal{G}(u^{n}),$ (B.2a)
$u^{n+1}$ $=(2-\sqrt{3})u^{n}-\frac{\sqrt{3}-1}{2}\Delta t\mathcal{G}(u^{n})+(\sqrt{3}-1)u^{(2)}+\Delta t\mathcal{F}(u^{(2)}).$ (B.2b)
However, with the scaling limiters, denoted by $\mathcal{L}$ or indicated by a tilde symbol, Scheme 1L in (B.3) and Scheme 2L in (B.4) are no longer equivalent.
Scheme 1L:
$u^{(2)}$ $=\tilde{u}^{n}+\frac{\Delta t}{2}\mathcal{G}(\tilde{u}^{n}),\quad$ $\tilde{u}^{(2)}$ $=\mathcal{L}({u}^{(2)}),$ (B.3a)
$u^{n+1}$ $=\tilde{u}^{n}+\Delta t\mathcal{F}(\tilde{u}^{(2)}),\quad$ $\tilde{u}^{n+1}$ $=\mathcal{L}({u}^{n+1}).$ (B.3b)
Scheme 2L:
$u^{(2)}$ $=\tilde{u}^{n}+\frac{\Delta t}{2}\mathcal{G}(\tilde{u}^{n}),\quad\tilde{u}^{(2)}=\mathcal{L}({u}^{(2)}),$ (B.4a)
$u^{n+1}$ $=(2-\sqrt{3})\tilde{u}^{n}-\frac{\sqrt{3}-1}{2}\Delta t\mathcal{G}(\tilde{u}^{n})+(\sqrt{3}-1)\tilde{u}^{(2)}+\Delta t\mathcal{F}(\tilde{u}^{(2)}),\quad\tilde{u}^{n+1}=\mathcal{L}({u}^{n+1}).$ (B.4b)
This is because, in Scheme 1L, (B.3b) can be rewritten as
$u^{n+1}=\tilde{u}^{n}+\Delta t\mathcal{F}(\tilde{u}^{(2)})=(2-\sqrt{3})\tilde{u}^{n}-\frac{\sqrt{3}-1}{2}\Delta t\mathcal{G}(\tilde{u}^{n})+(\sqrt{3}-1){u}^{(2)}+\Delta t\mathcal{F}(\tilde{u}^{(2)}).$ (B.5)
Comparing (B.4b) and (B.5), we note that (B.5) contains $u^{(2)}$ instead of $\tilde{u}^{(2)}$ in its third-term.
Note that Scheme 1L is the suggested implementation; however, it is the bound-preserving property of Scheme 2L that follows directly from Lemma 3.4.
Despite the difference between (B.5) and (B.4b), we can still prove that Scheme 1L is bound-preserving. Indeed, since ${u}^{(2)}$ and ${\tilde{u}}^{(2)}$ admit the same cell average, the cell averages of $u^{n+1}$ defined in (B.5) and (B.4b) are the same. To be more specific, for Scheme 1L, we have
$\overline{u}^{n+1}$ $=\overline{\tilde{u}^{n}+\Delta t\mathcal{F}(\tilde{u}^{(2)})}$ (B.6)
$=\overline{(2-\sqrt{3})\tilde{u}^{n}-\frac{\sqrt{3}-1}{2}\Delta t\mathcal{G}(\tilde{u}^{n})}+\overline{(\sqrt{3}-1){u}^{(2)}+\Delta t\mathcal{F}(\tilde{u}^{(2)})}$
$=\overline{(2-\sqrt{3})\tilde{u}^{n}-\frac{\sqrt{3}-1}{2}\Delta t\mathcal{G}(\tilde{u}^{n})}+\overline{(\sqrt{3}-1)\tilde{u}^{(2)}+\Delta t\mathcal{F}(\tilde{u}^{(2)})}$
$=(2-\sqrt{3})\left(\overline{\tilde{u}^{n}-\frac{\sqrt{3}+1}{2}\Delta t\mathcal{G}(\tilde{u}^{n})}\right)+(\sqrt{3}-1)\left(\overline{\tilde{u}^{(2)}+\frac{\sqrt{3}+1}{2}\Delta t\mathcal{F}(\tilde{u}^{(2)})}\right).$
By Lemma 3.4, we can conclude that Scheme 1L is still bound-preserving.
B.2 Proof of the general case in Theorem 3.5
Now we consider the general case in Theorem 3.5. With limiters, the cRKDG method can be written as
$u^{(i)}$ $=\tilde{u}^{n}+\Delta t\sum_{j=1}^{i-1}a_{ij}\mathcal{G}\left(\tilde{u}^{(j)}\right),\quad\tilde{u}^{(i)}=\mathcal{L}(u^{(i)}),\quad i=1,2,\cdots,s,$ (B.7a)
$u^{n+1}$ $=\tilde{u}^{n}+\Delta t\sum_{i=1}^{s}b_{i}\mathcal{F}\left(\tilde{u}^{(i)}\right),\quad\tilde{u}^{n+1}=\mathcal{L}(u^{n+1}).$ (B.7b)
We first show that all inner stages $\tilde{u}^{(i)}$ preserve the intended bound.
Proposition B.1.
Let $0\leq\beta_{ij}\leq 1$ and $\sum_{j=1}^{i-1}\beta_{ij}=1$. Suppose that for the scheme without limiters, or equivalently, when $\mathcal{L}=\mathcal{I}$ is the identity operator in (B.7), we have
$u^{(i)}=u^{n}+\Delta t\sum_{j=1}^{i-1}a_{ij}\mathcal{G}(u^{(j)})=\sum_{j=1}^{i-1}{\beta_{ij}}\left(u^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(u^{(j)})\right).$ (B.8)
Then for the scheme with limiters (B.7), we have
$u^{(i)}=\tilde{u}^{n}+\Delta t\sum_{j=1}^{i-1}a_{ij}\mathcal{G}(\tilde{u}^{(j)})=\sum_{j=1}^{i-1}{\beta_{ij}}\left(u^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(j)})\right).$ (B.9)
Let us admit Proposition B.1 and defer its proof to the next subsection.
By taking the cell average on both sides of (B.9) and noting the fact that the limiter doesn’t alter cell averages, namely, $\overline{u}^{(j)}=\overline{\tilde{u}}^{(j)}$, we have
$\overline{u}^{(i)}=\sum_{j=1}^{i-1}{\beta_{ij}}\left(\overline{u}^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\overline{\mathcal{G}(\tilde{u}^{(j)})}\right)=\sum_{j=1}^{i-1}{\beta_{ij}}\left(\overline{\tilde{u}^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(j)})}\right).$ (B.10)
Therefore, by Lemma 3.4, we have $\tilde{u}^{(i)}\in B$ under the prescribed conditions.
Now we prove that $\tilde{u}^{n+1}$ preserves the intended bound. For the final stage, one can similarly show that
$u^{n+1}$ $=u^{n}+\Delta t\sum_{i=1}^{s}b_{i}\mathcal{F}\left(\tilde{u}^{(i)}\right)$ (B.11)
$=\sum_{i=1}^{s}\beta_{i}^{\mathcal{F}}\left(u^{(i)}+\Delta t\mu_{i}^{\mathcal{F}}\mathcal{F}(\tilde{u}^{(i)})\right)+\sum_{i=1}^{s}\beta_{i}^{\pm\mathcal{G}}\left(u^{(i)}\pm\Delta t\mu_{i}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(i)})\right),$
if the scheme without limiters admits a similar convex decomposition. Here $0\leq\beta_{i}^{\mathcal{F}},\beta_{i}^{\pm\mathcal{G}}\leq 1$ and $\sum_{i=1}^{s}\beta_{i}^{\mathcal{F}}+\sum_{i=1}^{s}\beta_{i}^{\pm\mathcal{G}}=1$. After taking the cell average on both hand sides, we have
$\overline{u}^{n+1}$ $=\sum_{i=1}^{s}\beta_{i}^{\mathcal{F}}\left(\overline{u^{(i)}+\Delta t\mu_{i}^{\mathcal{F}}\mathcal{F}(\tilde{u}^{(i)})}\right)+\sum_{i=1}^{s}\beta_{i}^{\pm\mathcal{G}}\left(\overline{u^{(i)}\pm\Delta t\mu_{i}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(i)})}\right)$ (B.12)
$=\sum_{i=1}^{s}\beta_{i}^{\mathcal{F}}\left(\overline{\tilde{u}^{(i)}+\Delta t\mu_{i}^{\mathcal{F}}\mathcal{F}(\tilde{u}^{(i)})}\right)+\sum_{i=1}^{s}\beta_{i}^{\pm\mathcal{G}}\left(\overline{\tilde{u}^{(i)}\pm\Delta t\mu_{i}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(i)})}\right).$
Once again, by Lemma 3.4, we have $\tilde{u}^{n+1}\in B$ under the prescribed conditions.
B.3 Proof of Proposition B.1
The key to the proof is the following. To derived (B.8), we have to use the identity
$u^{(j)}=u^{n}+\Delta t\sum_{\ell=1}^{j-1}a_{j\ell}\mathcal{G}(u^{(\ell)}).$ (B.13)
But in the scheme with limiters, this relation is replaced by
$u^{(j)}=\tilde{u}^{n}+\Delta t\sum_{\ell=1}^{j-1}a_{j\ell}\mathcal{G}(\tilde{u}^{(\ell)}).$ (B.14)
Thus we will have $u^{(j)}$ instead of $\tilde{u}^{(j)}$ for the first term inside the parenthesis on the right hand side of (B.9).
The detailed verification is given as follows. For the scheme without limiters, with (B.13) and the fact $\sum_{j=1}^{i-1}\beta_{ij}=1$, we have
$\sum_{j=1}^{i-1}{\beta_{ij}}\left(u^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(u^{(j)})\right)$ $=\sum_{j=1}^{i-1}{\beta_{ij}}\left(u^{n}+\Delta t\sum_{\ell=1}^{j-1}a_{j\ell}\mathcal{G}(u^{(\ell)})\right)\pm\Delta t\sum_{j=1}^{i-1}{\beta_{ij}}\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(u^{(j)})$ (B.15)
$=u^{n}+\Delta t\sum_{\ell=1}^{i-1}\sum_{j=1}^{\ell-1}{\beta_{i\ell}}a_{\ell j}\mathcal{G}(u^{(j)})\pm\Delta t\sum_{j=1}^{i-1}{\beta_{ij}}\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(u^{(j)})$
$=u^{n}+\Delta t\sum_{j=1}^{i-2}\sum_{\ell=j+1}^{i-1}{\beta_{i\ell}}a_{\ell j}\mathcal{G}(u^{(j)})\pm\Delta t\sum_{j=1}^{i-1}{\beta_{ij}}\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(u^{(j)})$
$=u^{n}+\Delta t\sum_{j=1}^{i-1}\left(\pm\beta_{ij}\mu_{ij}^{\pm\mathcal{G}}+\sum_{\ell=j+1}^{i-1}{\beta_{i\ell}}a_{\ell j}\right)\mathcal{G}(u^{(j)}).$
Here we have used the fact $a_{\ell,i-1}=0$ for $j+1\leq\ell\leq i-1$ in the last equality. Therefore, comparing both hand sides of the second equality in (B.8), we have
$a_{ij}=\pm\beta_{ij}\mu_{ij}^{\pm\mathcal{G}}+\sum_{\ell=1}^{i-1}{\beta_{i\ell}}a_{\ell j}.$ (B.16)
As a result, for the scheme with limiters, by reversing the derivation in (B.15) and applying (B.14) in the end, we have
$u^{(i)}$ $=\tilde{u}^{n}+\Delta t\sum_{j=1}^{i-1}a_{ij}\mathcal{G}(\tilde{u}^{(j)})$ (B.17)
$=\tilde{u}^{n}+\Delta t\sum_{j=1}^{i-1}\left(\pm\beta_{ij}\mu_{ij}^{\pm\mathcal{G}}+\sum_{\ell=j+1}^{i-1}{\beta_{i\ell}}a_{\ell j}\right)\mathcal{G}(\tilde{u}^{(j)})$
$=\sum_{j=1}^{i-1}{\beta_{ij}}\left(\left(\tilde{u}^{n}+\Delta t\sum_{\ell=1}^{j-1}a_{j\ell}\mathcal{G}(\tilde{u}^{(\ell)})\right)\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(j)})\right)$
$=\sum_{j=1}^{i-1}{\beta_{ij}}\left(u^{(j)}\pm\Delta t\mu_{ij}^{\pm\mathcal{G}}\mathcal{G}(\tilde{u}^{(j)})\right).$
This completes the proof of Proposition B.1.